OpenStudy (anonymous):

Watch like a 1000 people come to this.

OpenStudy (anonymous):

Saying please close this question

OpenStudy (anonymous):

Do you have a question? Or are you just spamming?

OpenStudy (anonymous):

Im trying to find help but noone helps.

OpenStudy (anonymous):

What do you need help WITH?

OpenStudy (anonymous):

Use a transformation to solve the equation. z – (–0.3) = –2.4 A. z = –2.7 B. z = –2.1 C. z = 2.1 D. z = 2.7

OpenStudy (anonymous):

I got A...

OpenStudy (anonymous):

Thats what i got.

OpenStudy (anonymous):

Just making sure.

OpenStudy (anonymous):

z - (-0.3) = -2.4 z + 0.3 = -2.4 z = -2.4 - 0.3 z = -2.4 + -0.3 z = -2.7